OpenStudy (anonymous):

234.51ha change to m

OpenStudy (anonymous):

Ha is used to quantify area. m is used to quantify distances. I think u typed wrongly? It should be m^2 or square meters. Anw 1 ha is equivalent to 10 000 square meters. So just multiply to get the answer.

OpenStudy (anonymous):

calculate the circumference,area and radius of the circle

OpenStudy (anonymous):

I assume the striaght line u draw in the circle is 9cm. That is the diameter. Given diameter, the radius is 9/2= 4.5cm (note diameter is 2 times of radius) Circumference is given by formula= 2*pi*r (note r stands for radius) = 2*pi*4.5 =28.3cm Area of circle is given by formula = pi*r^2 = pi* 4.5^2 =63.6 cm^2